### v1

@@ Abstract
As the Large-Language-Model-driven (LLM-driven) Artificial Intelligence (AI) bots became popular, people realized their strong potential in Task-Oriented Dialogue (TOD). However, bots relying wholly on LLMs are unreliable in their knowledge, and whether they can finally produce a correct result for the task is not guaranteed. The collaboration among these agents also remains a challenge, since the necessary information to convey is unclear, and the information transfer is by prompts — unreliable, and malicious knowledge is easy to inject. With the help of logic programming tools such as Answer Set Programming (ASP), conversational agents can be built safely and reliably, and communication among the agents made more efficient and secure. We proposed an Administrator-Assistant Dual-Agent paradigm, where the two ASP-driven bots share the same knowledge base and complete their tasks independently, while the information can be passed by a Collaborative Rule Set (CRS). The knowledge and information conveyed are encapsulated and invisible to the users, ensuring the security of information transmission. We have constructed AutoManager, a dual-agent system for managing the drive-through window of a fast-food restaurant such as Taco Bell in the US. In AutoManager, the assistant bot takes the customer’s order while the administrator bot manages the menu and food supply. We evaluated our AutoManager and compared it with the real-world Taco Bell Drive-Thru AI Order Taker, and the results show that our method is more reliable.
keywords:
Answer Set Programming; s(CASP); Large Language Model; chatbot
1 Introduction
@@
In our earlier work, autoconcierge; autocompanion have developed reliable task-oriented chatbots as well as socialbots that interact with humans. We used a large language model as a semantic parser and answer set programming (ASP) for performing the backend reasoning. The LLM-based semantic parser turns text into knowledge represented as logic predicates. ASP is then used for performing reasoning using these predicates. Pre-existing commonsense knowledge represented in ASP is also used for this purpose. The computed response is a set of predicates that are then converted into text, using the LLM in reverse.
These chatbots operate as a single agent that interacts with a human user. In this paper, we report on the development of collaborative chatbots that are reliable. We limit ourselves to the dual-agent collaboration in this paper. That is, only two chatbots working collaboratively are considered. It consists of an administrator bot that can modify the knowledge base and an assistant bot for serving the customer, so that a human manager can change the parameters of the assistant bot’s conversation with other humans. Each collaborative chatbot performs a different but complementary task, and communicates with each other through a shared knowledge base. Each chatbot interacts with a different class of human users.
@@
Our motivation for the dual-agent chatbot system is the automated management of a fast-food restaurant, such as McDonald’s or Taco Bell. In such a fast-food restaurant, a person at the drive-through window takes the order after interacting with the customer, while a manager inside the restaurant keeps track of the resources needed to fulfill the order. The manager may instruct the person taking orders not to take orders for fries anymore, for example, as the restaurant has run out of fries. Based on the dual-agent collaboration system, we built AutoManager to model the collaboration for running a Taco Bell restaurant, where the administrator bot manages the food supply and menu updates, and the assistant bot takes the customers’ orders. The assistant bot will talk to the customers, while the manager bot will talk to an actual manager who will provide instructions in natural language. We evaluated the conversation quality in the AutoManager manually and compared it with the real-world Taco Bell Drive-Thru AI system, which is largely based on machine learning technology. The result shows that our AutoManager, based on a hybrid of LLM and ASP, is more reliable than the real-world Taco Bell AI.
The main contribution of our research is to show how a combination of LLM and ASP can be used to develop collaborative bots. We developed a general methodology for developing such collaborative bots, where each bot has a different function and interacts with a certain class of users. We illustrate our ideas by constructing a dual-agent bot for managing a Taco Bell restaurant. Notably, recent attempts to develop similar chatbots by large companies have failed because they primarily relied on machine learning, reported by failed-mcdonald-bot. Existing solutions that have had moderate success by tacobell-bot also need considerable customized training and frequent error-fixing from human beings. Therefore, they are essentially assistants to humans rather than systems that can operate independently. Again, over-reliance on machine learning technology is the main culprit. Hence, our research also shows how automated reasoning based on answer set programming leads to developing reliable and robust AI systems.
2 Background
2.1 Large Language Models (LLMs):
@@
Until recently, transformer-based Large Language Models, pre-trained on an enormous quantity of well-annotated data, have been applied to general NLP tasks. With the advent of Large Language Models, the paradigm changed from pre-training and fine-tuning (by pre_trained_transformers) to teaching a language model any arbitrary task using just a few demonstrations, called in-context learning, or prompt engineering. gpt3 introduced an LLM called GPT-3 containing approximately 175 billion parameters that have been trained on a massive corpus of filtered online text, on which the well-known ChatGPT is based. GPT-3, as well as the later GPT-4 (by gpt4), can perform competitively on several tasks such as question-answering, semantic parsing, and machine translation. However, such LLMs lack the ability of precise logical reasoning and find it hard to overcome the hallucination brought about by the training data, according to gpt3-scope; chain; chatgpt-critic.
2.2 Answer Set Programming and s(CASP)
@@
Answer Set Programming (ASP), according to cacm-asp and gupta-csr, is a logic programming paradigm suited for knowledge representation and reasoning that facilitates commonsense reasoning. The s(CASP) system proposed by scasp is an answer set programming system that supports predicates, constraints over non-ground variables, uninterpreted functions, and, most importantly, a top-down, query-driven execution strategy. These features make it possible to return answers with non-ground variables (possibly including constraints among them) and compute partial models by returning only the fragment of a stable model necessary to support the answer.
@@
Complex commonsense knowledge can be represented in ASP, and the s(CASP) query-driven predicate ASP system can be used for querying it. Commonsense knowledge can be emulated using (i) default rules, (ii) integrity constraints, and (iii) multiple possible worlds by gelfondkahl; gupta-csr. Default rules are used for jumping to a conclusion in the absence of exceptions, e.g., a bird normally flies, unless it’s a penguin.
@@
A large number of commonsense reasoning applications have already been developed using ASP and the s(CASP) system by blawx; logical-english; chef; murder-trial. The query-driven s(CASP) ASP system was crucial for the AutoConcierge system developed by autoconcierge earlier. s(CASP) was used to perform commonsense reasoning resembling a human concierge: (i) To check for consistency and completeness in the knowledge derived from the user’s dialogs, and to ask further questions and seek more information in case it was inconsistent or incomplete. (ii) Subsequently, concluding the extracted knowledge. Explanation for each response can also be given by justifications for successful queries as proof trees, according to scasp-justification. autocompanion also developed a Socialbot that can hold conversations with a stranger about movies, books, and sports using similar techniques.
2.3 The STAR Framework
@@
This mechanism is embodied in the Semantic-parsing Transformer and ASP Reasoner (STAR) framework, proposed by star, leveraging the advantages of LLMs and ASP systems by combining them systematically. It parses the semantics of the text sentences to generate the predicates using LLMs such as those from the GPT series. After that, it sends the predicates to the ASP system to get reliable answers through reasoning. This framework was previously applied to task-oriented chatbots and socialbots, such as AutoConcierge by autoconcierge and AutoCompanion by autocompanion, and has proved to work well.
3 Dual-Agent Chatbot Design
@@
The dual-agent collaborative system, as the name indicates, consists of two conversational agents that share the knowledge base but have different functions. One typical use of the dual system is the administrator-assistant system, where the administrator manages and maintains the knowledge base information, while the assistant offers services to the customers using the knowledge base. In a restaurant, for example, the administrator maintains the menu and food supplies, and the assistant takes orders and serves customers.
@@
The administrator and assistant bot are both constructed based on the STAR framework, in which the LLM is only used for parsing natural language into a pre-defined predicate logic-based vocabulary and vice versa. All the decision-making and consistency-checking steps are done by a reasoning engine driven by the s(CASP) goal-directed ASP system from scasp. This framework ensures the reliability of the decision and the robustness of the input and output. The administrator and assistant bots’ functions in AutoManager are shown in Figure 1 and will be detailed below.
@@
Figure 1: AutoManager Architecture. The administrator and assistant bot are constructed in the STAR framework. Two bots share the knowledge base, temporary information list, and a collaborative rule set.
3.1 Administrator Bot Functions
@@
In AutoManager, the administrator maintains the menu and food supplements. The users (who represent managers) can change the menu by adding, deleting, or modifying the menu items, and notifying the shortage or restoration of some ingredients. In this work, we use “dish” to represent any item that can be ordered on the menu except combos, ingredients, and sauces. On the other hand, the term “food” refers to the general name for the combo, dish, ingredient, and sauce.
@@
LLM Parser and Ontology. In AutoManager, both the administrator and assistant bot leverage an LLM (GPT-4) for extracting knowledge from a manager’s and customer’s dialog respectively. This knowledge is represented as predicates, that come from a pre-determined vocabulary. Per the STAR framework, this logical vocabulary can be thought of as a language that the bot “understands”. The LLM is used primarily to translate natural language to logic predicates and vice versa. The predicates corresponding to a natural language sentence can be thought of as representing its meaning or linguistic deep structure.
@@
To capture the main functions of the administrator, the following predicates have been pre-defined: runout, restore, add, edit, and delete. These predicates have distinct meanings wrt the commonsense reasoning rules defined for the administrator bot. The predicate runout/1 and restore/1 indicate that a certain ingredient is out of stock or has been restored. The predicate add(Type, Food) means to add a kind of food to the menu, where the first arity is the type to add, and the second arity is the name of the food. This kind of food can either be a dish (e.g., Grilled Beef Taco), a combo of dishes (e.g., Grilled Beef Taco Combo), or an ingredient (e.g., Grilled Beef). If the user does not specify the type of food, the LLM parser will extract add(Food) instead. Otherwise, “Type” can be either “dish”, “combo”, or “ingredient”. If the user wants to delete a food from the menu, the predicate delete(Food) is what they need. The predicates add(Food, Property, Value), edit(Food, Property, Value), edit(Food, Property, Original_Value, New_Value), delete(Food, Property, Value), and delete(Food, Property) express the users’ intentions of modifying the property of some food to some value. For example, if the manager wants to change the price of the Grilled Cheese Burrito to \$3.80, the predicate generated would be: edit(‘Grilled Cheese Burrito’, price, 3.80). If the manager wants to change one of the ingredients of Crunchy Taco from cheese to lettuce, the predicate edit(‘Crunchy Taco’, included_ingredient, ‘Cheese’, ‘Lettuce’). will be generated. Also, predicate quit/0 and irrelevant/0 show that the user wants to leave, or said something irrelevant. These predicates, along with their usage, are sent as prompts to the LLM, where the prompt also includes a few simple examples to define the format that the LLM will follow.
@@
Conversational Knowledge Template (CKT). CKTs, proposed by ckt, are state machines that represent the logic that a bot must follow to generate a response. The CKTs are coded in ASP. One characteristic of human input is that it can contain variable information. For example, a user may say “I want to order a soft taco,” and then say “Can you please replace meat with beans?” or a user may say “I want to order a soft taco with beans instead of meat.” CKTs can handle a variable amount of information contained in each sentence. In the administrator bot, to add a type of food, CKT will loop in asking for the information listed below: the category of the food, its price, its ingredients, the calories it contains, and the popular toppings to recommend. As for adding a combo, CKT will ask for its price, calories contained, and the food it includes. Once the CKT decides what to ask for in the next round of dialogue, a predicate ask(Name, Property) will be generated. Note that CKTs are flexible structures. The natural language input can be provided by a human in one shot or piecemeal over multiple responses.
@@
Response Generation. The response generated by the administrator bot includes two parts: confirmation and question. The predicates for confirmation are directly correspond to the user’s input. For instance, when the user says that the lettuce is out of stock, predicate confirm(‘run out’, ‘Lettuce’) will be applied as the confirmation, which will later be sent to an LLM for generating a natural language sentence like “We are short of lettuce, understood.” The question part is only for the case where a CKT is invoked, e.g., adding a new dish to the menu. Similar to the predicate extraction part, we provide the LLM with the usage of each predicate and several examples in the prompt. We also add instructions to the prompt to tell the LLM the background of the dialogue; consequently, the LLM can generate suitable sentences as responses.
3.2 Assistant Bot Functions

### v2

@@ Abstract
As the Large-Language-Model-driven (LLM-driven) Artificial Intelligence (AI) bots became popular, people realized their strong potential in Task-Oriented Dialogue (TOD). However, bots relying wholly on LLMs are unreliable in their knowledge, and whether they can finally produce a correct outcome for the task is not guaranteed. The collaboration among these agents also remains a challenge, since the necessary information to convey is unclear, and the information transfer is by prompts—unreliable, and malicious knowledge is easy to inject. With the help of knowledge representation and reasoning tools such as Answer Set Programming (ASP), conversational agents can be built safely and reliably, and communication among the agents made more reliable as well. We propose an Administrator-Assistant Dual-Agent paradigm, where ASP-driven bots share the same knowledge base and complete their assigned tasks independently. The agents communicate with each other through the knowledgebase ensuring consistency. The knowledge and information conveyed are encapsulated and invisible to the users, ensuring the security of information transmission. To illustrate the dual-agent conversational agent paradigm, we have constructed AutoManager, a dual-agent system for managing the drive-through window of a fast-food restaurant such as Taco Bell in the US. In AutoManager, the assistant bot takes the customer’s order while the administrator bot manages the menu and food supply. We evaluated our AutoManager system and compared it with the real-world Taco Bell Drive-Thru AI Order Taker, and the results show that our method is more reliable.
Index Terms:
Answer Set Programming, s(CASP), Large Language Model, chatbot
I Introduction
Collaboration is ubiquitous in human society and can even be regarded as the very foundation upon which societies and civilizations are built. Enterprises, for example, require diverse roles to fulfill specialized functions—marketing, technology, finance, legal affairs, etc.—and only through collaboration among these distinct positions can a company operate effectively. The same principle applies to conversational agents. Different agents serve varied user groups and accomplish distinct tasks, while their inter-agent collaboration sustains the harmonious operation of the entire system.
For instance, consider automating an airport check-in operation with AI agents. An Operations Management bot could allocate boarding gates and respond to flight delays or emergencies, while a Counter Assistant bot assists passengers with check-in, seat selection, baggage check-in, class upgrades, and real-time information dissemination. Things may, however, change dynamically on the ground, e.g., an airport gate suffers a mechanical fault and the aircraft has to be re-assigned to another gate. In such scenarios, the real-time and accurate information exchange between agents and the dynamic and contextually appropriate responses to received information demands heightened attention.
Collaboration between such AI bots requires information known to one bot to be made instantaneously available to the other. For conversational bots solely based on large language models (LLMs), the instantaneous information can only be passed through prompts and cannot be obtained through training or fine-tuning (though Therefore, instantaneous information will occupy a large amount of prompt space, thereby reducing the operating efficiency and reliability of the LLM. In addition, due to the unreliable reasoning and hallucinatory nature of LLMs, the instantaneous information cannot always be used properly to complete the tasks. Therefore, task-oriented bots that are based on LLMs have not always succeeded. A recent unsuccessful attempt is reported by McDonald [1] for developing drive-through window order-taking chatbots. A common issue arose from the absence of a reliable communication mechanism: the AI order taker fails to receive real-time updates on ingredient availability in restaurants, for example. Even when such information is accurately transmitted, the chatbot may neglect or lack sufficient reasoning capacity to initiate appropriate actions, such as notifying customers of shortages, resulting in the inability to fill an order, for example. Another AI order taker, delployed by Taco Bell [2], also needs considerable customized training along with constant supervision by a human being. Therefore, they are essentially assistants to humans rather than systems that can operate independently.
Motivated by the above problem, we propose the development of reliable knowledge-driven collaborative chatbots. Different from other agentic AI systems, our method focuses on the accuracy and controllability of the agent-wide communication, and employs the reasoning system to make the communication reliable. We limit ourselves to the dual-agent collaboration in this paper, i.e., only two chatbots working collaboratively are considered. Typically, such a dual-agent system consists of a manager bot to whom changes brought about by instantaneous events are communicated by the human manager, and a customer service chatbot that helps human users (human customers).
@@ I Introduction
In our earlier work, we developed task-oriented chatbots [3] as well as socialbots [4] that are knowledge-driven and thus reliable. To ensure the reliability of chatbot responses, we limited the use of LLMs as a semantic parser. We used answer set programming (ASP), a knowledge representation and reasoning formalism, to perform the backend reasoning. The reasoning that we perform is true deductive or abductive reasoning and is distinct from “reasoning” performed by Large Reasoning Models. The LLM-based semantic parser turns text into knowledge represented as logic predicates. ASP is then used for performing reasoning over these predicates. Pre-existing commonsense knowledge represented in ASP is also employed during this reasoning. The computed response is a set of predicates that are then converted into text, using an LLM as a reverse logic predicates to English text translator. Each chatbot in the collaboration system is constructed following the above method. Each single bot performs a distinct but complementary task by interacting with a specific type of human user, and communicates with the other through a shared knowledge base.
@@
Based on the dual-agent collaboration system, we built AutoManager to model the collaboration for running a Taco Bell restaurant, where the manager bot manages the food supply and menu updates, and the customer service bot takes customer orders. The customer service bot chats with the customers to directly take orders, while the manager bot chats with an actual human manager who provides instructions in natural language to keep the shared knowledge consistent. We evaluated the conversation quality in the AutoManager manually and compared it with the real-world Taco Bell Drive-Thru AI system, which is largely based on machine learning technology. The result shows that our AutoManager, based on a hybrid of LLM and ASP, is more reliable than the real-world Taco Bell AI.
The main contribution of our paper is the KRR-driven multi-agent collaborating chatbot paradigm and its use in developing a practical reliable application that was hitherto not possible.
II Background
II-A Large Language Models (LLMs):
@@
Until recently, transformer-based Large Language Models, pre-trained on an enormous quantity of well-annotated data, have been applied to general NLP tasks. With the advent of Large Language Models, the paradigm changed from pre-training and fine-tuning (by [5]) to teaching a language model any arbitrary task using just a few demonstrations, called in-context learning, or prompt engineering. [6] introduced an LLM called GPT-3 containing approximately 175 billion parameters that have been trained on a massive corpus of filtered online text, on which the well-known ChatGPT is based. GPT-3, as well as the later GPT-4 (by [7]), can perform competitively on several tasks such as question-answering, semantic parsing, and machine translation. However, such LLMs lack the ability of precise logical reasoning and find it hard to overcome the hallucination brought about by the training data, according to [8, 9, 10].
II-B Answer Set Programming and s(CASP)
@@
Answer Set Programming (ASP) [11, 12] is a logic programming paradigm suited for knowledge representation and reasoning that facilitates commonsense reasoning. The s(CASP) system [13] is an answer set programming system that supports predicates, constraints over non-ground variables, uninterpreted functions, and, most importantly, a top-down, query-driven execution strategy. These features make it possible to return answers with non-ground variables (possibly including constraints among them) and compute partial models by returning only the fragment of a stable model necessary to support the answer.
@@
Complex commonsense knowledge can be represented in ASP, and the s(CASP) query-driven predicate ASP system can be used for querying it. Commonsense knowledge can be emulated using (i) default rules, (ii) integrity constraints, and (iii) multiple possible worlds by [14, 12]. Default rules are used for jumping to a conclusion in the absence of exceptions, e.g., a bird normally flies, unless it’s a penguin.
@@
A large number of commonsense reasoning applications have already been developed using ASP and the s(CASP) system by [15, 16, 17, 18]. The query-driven s(CASP) ASP system was crucial for the AutoConcierge system developed by [3] earlier. s(CASP) was used to perform commonsense reasoning resembling a human concierge: (i) To check for consistency and completeness in the knowledge derived from the user’s dialogs, and to ask further questions and seek more information in case it was inconsistent or incomplete. (ii) Subsequently, concluding the extracted knowledge. Explanation for each response can also be given by justifications for successful queries as proof trees, according to [19]. [4] also developed a Socialbot that can hold conversations with a stranger about movies, books, and sports using similar techniques.
II-C The STAR Framework
@@
This mechanism is embodied in the Semantic-parsing Transformer and ASP Reasoner (STAR) framework, proposed by [20], leveraging the advantages of LLMs and ASP systems by combining them systematically. It parses the semantics of the text sentences to generate the predicates using LLMs such as those from the GPT series. After that, it sends the predicates to the ASP system to get reliable answers through reasoning. This framework was previously applied to task-oriented chatbots and socialbots, such as AutoConcierge by [3] and AutoCompanion by [4], and has proved to work well.
III Dual-Agent Chatbot Design
@@
As the name indicates, the dual-agent collaborative system consists of two conversational agents who share the knowledge base but have different functions. In such a dual system, the manager bot manages and maintains the knowledge base information, while the customer service bot offers services to customers using the knowledge base. Typically, in a restaurant, for example, the manager bot maintains the menu and food supplies, and the customer service bot takes orders and answers the customers’ questions.
@@
Both bots are constructed based on the STAR framework, in which the LLM is only used for parsing natural language into a pre-defined predicate logic-based vocabulary and vice versa. All the decision-making and consistency-checking steps are done by a reasoning engine driven by the s(CASP) goal-directed ASP system from [13]. This framework ensures the reliability of the decision and the robustness of the input and output. The administrator and assistant bots’ functions in AutoManager are shown in Figure 1 and will be detailed below.
@@
Figure 1: AutoManager Architecture. The manager and customer service bot are constructed in the STAR framework. Two bots share the knowledge base, temporary information list, and a collaborative rule set.
III-A Manager Bot Functions
@@
In AutoManager, the manager bot maintains the menu and food supplements. The users (who represent restaurant employees and managers) can change the menu by adding, deleting, or modifying the menu items, and notifying of the shortage or restoration of some ingredients. In this work, we use “dish” to represent any item that can be ordered on the menu except combos, ingredients, and sauces. On the other hand, the term “food” refers to the general name for the combo, dish, ingredient, and sauce.
@@
LLM Parser and Ontology. In AutoManager, both the manager and customer service bot leverage an LLM (GPT-4) for extracting knowledge from a manager’s and customer’s dialog respectively. This knowledge is represented as predicates, that come from a pre-determined vocabulary. Per the STAR framework, this logical vocabulary can be thought of as a language that the bot “understands”. The LLM is used primarily to translate natural language to logic predicates and vice versa. The predicates corresponding to a natural language sentence can be thought of as representing its meaning or linguistic deep structure.
@@
To capture the main functions of the manager bot, the following predicates have been pre-defined: runout, restore, add, edit, and delete. These predicates have distinct meanings wrt the commonsense reasoning rules defined for the administrator bot. The predicate runout/1 and restore/1 indicate that a certain ingredient is out of stock or has been restored. The predicate add(Type, Food) means to add a kind of food to the menu, where the first arity is the type to add, and the second arity is the name of the food. This kind of food can either be a dish (e.g., Grilled Beef Taco), a combo of dishes (e.g., Grilled Beef Taco Combo), or an ingredient (e.g., Grilled Beef). If the user does not specify the type of food, the LLM parser will extract add(Food) instead. Otherwise, “Type” can be either “dish”, “combo”, or “ingredient”. If the user wants to delete a food from the menu, the predicate delete(Food) is what they need. The predicates add(Food, Property, Value), edit(Food, Property, Value), edit(Food, Property, Original_Value, New_Value), delete(Food, Property, Value), and delete(Food, Property) express the users’ intentions of modifying the property of some food to some value. For example, if the restaurant runner wants to change the price of the Grilled Cheese Burrito to \$3.80, the predicate generated would be: edit(‘Grilled Cheese Burrito’, price, 3.80). And if the runner wants to change one of the ingredients of Crunchy Taco from cheese to lettuce, the predicate edit(‘Crunchy Taco’, included_ingredient, ‘Cheese’, ‘Lettuce’). will be generated. Also, predicate quit/0 and irrelevant/0 show that the user wants to leave, or said something irrelevant. These predicates, along with their usage, are sent as prompts to the LLM, where the prompt also includes a few simple examples to define the format that the LLM will follow.
@@
Conversational Knowledge Template (CKT). CKTs, proposed by [21], are state machines that represent the logic that a bot must follow to generate a response. The CKTs are coded in ASP. One characteristic of human input is that it can contain variable information. For example, a user may say “I want to order a soft taco,” and then say “Can you please replace meat with beans?” or a user may say “I want to order a soft taco with beans instead of meat.” CKTs can handle a variable amount of information contained in each sentence. In the manager bot, to add a type of food, CKT will loop in asking for the information listed below: the category of the food, its price, its ingredients, the calories it contains, and the popular toppings to recommend. As for adding a combo, CKT will ask for its price, calories contained, and the food it includes. Once the CKT decides what to ask for in the next round of dialogue, a predicate ask(Name, Property) will be generated. Note that CKTs are flexible structures. The natural language input can be provided by a human in one shot or piecemeal over multiple responses.
@@
Response Generation. The response generated by the manager bot includes two parts: confirmation and question. The predicates for confirmation are directly correspond to the user’s input. For instance, when the user says that the lettuce is out of stock, predicate confirm(‘run out’, ‘Lettuce’) will be applied as the confirmation, which will later be sent to an LLM for generating a natural language sentence like “We are short of lettuce, understood.” The question part is only for the case where a CKT is invoked, e.g., adding a new dish to the menu. Similar to the predicate extraction part, we provide the LLM with the usage of each predicate and several examples in the prompt. We also add instructions to the prompt to tell the LLM the background of the dialogue; consequently, the LLM can generate suitable sentences as responses.
III-B Customer Service Bot Functions